# Quillen’s conjecture and unitary groups

### v3

@@ 2.4. Symmetric group and certain subsets of it
The minimal cases $n=2,3$ are consequence of [3, Proposition 4.8], as we are excluding $(p,q)=(3,2)$, and $(p,q)=(3,8)$ if $B$ contains field automorphisms of order $p$. Thus, in the rest of the proof, we assume that $n\geq 4$. As additional preparation, we refer the reader to the description, in Proposition 2.2, of the $p$-extensions LBLB of $L=\operatorname{PSU}\nolimits_{n}(q)$ and their $p$-ranks, as well as the $p$-ranks of $\operatorname{PGU}\nolimits_{n}(q)$ and its extension by field automorphisms of order $p$. For conciseness, we explicitly consider the possible generators of $B$: an inner-diagonal automorphism of order $p$, $c_{[g]}$, and a field automorphism of order $p$, $c_{[h]}\Phi c_{[h^{-1}]}$, where $g,h\in\operatorname{GU}\nolimits_{n}(q)$, see Subsection 2.2. After conjugating $\operatorname{PSU}\nolimits_{n}(q)$ by $h^{-1}]$, we may assume that,
@@
(i) $B=1$: Then $LB=L$ and we note that Theorem 4.3 also proves that
$\widetilde{H}_{n-1}(|\mathcal{A}_{p}(\operatorname{GU}\nolimits_{n}(q))_{>Z_{p}}|;\mathbb{Z})\neq 0,$
where $Z_{p}$ is the order-$p$ part of the center $Z$ of $\operatorname{GU}\nolimits_{n}(q)$ and where we exclude the case $(p,q)=(3,2)$. Moreover, if $n_{p}<(q+1)_{p}$, we use the inclusion of posets in [3, Lemma 4.7(a)],
$\mathcal{A}_{p}(\operatorname{GU}\nolimits_{n}(q))_{>Z_{p}}\subseteq\mathcal{A}_{p}(\operatorname{PSU}\nolimits_{n}(q)),$
where the groups involved have $p$-rank n1n-1. If $n_{p}\geq(q+1)_{p}$, then $p\mid n$ and, by [3, Lemma 4.7(b)], we get
@@
where the groups involved have $p$-rank n2n-2. Then, we can apply the case above, $n_{p}<(q+1)_{p}$, to the left-hand side as $p\nmid(n-1)$.
(ii) $B$ is cyclic generated by a field automorphism of order $p$: This case is similar to the previous one: Theorem 4.7 also proves that
$\widetilde{H}_{n-1}(|\mathcal{A}_{p}(\operatorname{GU}\nolimits_{n}(q)B)_{>Z_{p}}|;\mathbb{Z})\neq 0,$
where $Z_{p}$ is as in the previous case and where we exclude the case $(p,q)=(3,8)$. If $n_{p}<(q+1)_{p}$, we employ [3, Lemma 4.7(a)] after the addition of field automorphisms,
$\mathcal{A}_{p}(\operatorname{GU}\nolimits_{n}(q)B)_{>Z_{p}}\subseteq\mathcal{A}_{p}(\operatorname{PSU}\nolimits_{n}(q)B),$
where the groups involved have $p$-rank $n$. If $n_{p}\geq(q+1)_{p}$, we use [3, Lemma 4.7(b)] upon the addition of field automorphisms, and get
@@
where the groups involved have $p$-rank n1n-1. Note that $B$ acts on $\operatorname{PSU}\nolimits_{n-1}(q)$ by Equation (20). Again we may apply here the case $n_{p}<(q+1)_{p}$ of this point (ii) to the left-hand side as $p\nmid(n-1)$.
(iii) $B$ is cyclic generated by a diagonal automorphisms of order $p$: Equation (3) gives $\mathcal{A}_{p}(\operatorname{PSU}\nolimits_{n}(q)B)=\mathcal{A}_{p}(\operatorname{PGU}\nolimits_{n}(q))$ and this case follows from Theorem 4.3.
(iv) $B$ has $p$-rank 22 and it is generated by a field and a diagonal automorphism of order $p$: If $n_{p}<(q+1)_{p}$, we employ Theorem 4.7 and [3, Lemma 4.7(a)] after the addition of the field automorphism of order $p$, $\Phi$,
$\mathcal{A}_{p}(\operatorname{GU}\nolimits_{n}(q)\langle\Phi\rangle)_{>Z_{p}}\subseteq\mathcal{A}_{p}(\operatorname{PSU}\nolimits_{n}(q)\langle\Phi\rangle)\subseteq\mathcal{A}_{p}(\operatorname{PSU}\nolimits_{n}(q)B)\subseteq\mathcal{A}_{p}(\operatorname{PGU}\nolimits_{n}(q)B),$
and note that the groups involved have $p$-rank $n$. If $n_{p}\geq(q+1)_{p}$, we will employ [3, Lemma 4.7(b)] upon the addition of field automorphisms, but we need first the following digression: Write $q=s^{pl}$ and note that, by Equation (20), we have $B=\langle c_{[g]},\Phi\rangle$, where $c_{[g]}$ has already been conjugated by $h^{-1}]$. Now, let $\nu\in\mathbb{F}^{*}_{q^{2}}$ be an element of order $p$ with $\nu^{s^{l}+1}=1$ and consider the diagonal element of $\operatorname{GU}\nolimits_{n}(q)$ of order $p$,
$t=\operatorname{diag}(\nu,1,\ldots,1).$
Then $\Phi([t])=[t]$, $t]\in\operatorname{PGU}\nolimits_{n}(q)\setminus\operatorname{PSU}\nolimits_{n}(q)$ and, as the quotient group of Equation (3) is cyclic, there exists $s\in\operatorname{SU}\nolimits_{n}(q)$ with $[t]=[sg]$. Thus, $t]\in LB$ and we may assume that $B=\langle c_{[t]},\Phi\rangle$. Now, by [3, Proposition 4.7(b)], we have
where both groups have $p$-rank $n$. Then, we may apply the case $n_{p}<(q+1)_{p}$ of this point (iv) to the left-hand side, and we note that $B$ acts on $\operatorname{PSU}\nolimits_{n-1}(q)$ by the choice of the generators of $B$ we just made. ∎

### v5

@@ 2.4. Symmetric group and certain subsets of it
As preparation, we refer the reader to the account, in Proposition 2.2, of the $p$-extensions LBLB of $L=\operatorname{PSU}\nolimits_{n}(q)$ and their $p$-ranks, as well as the $p$-ranks of $\operatorname{PGU}\nolimits_{n}(q)$ and its extension by field automorphisms of order $p$.
As additional setup, we make a short detour to reformulate the conclusions of Theorems 4.3 and 4.7 in some cases: An a posteriori analysis of the cycles constructed in these two theorems, reveals that the elementary abelian $p$-subgroups appearing in the simplices involved in those cycles, have generators of the form
(19)  $\text{ either }{}^{{[g]}\!}{[}e]\text{ or }\Phi\text{ or }{}^{{[d]}\!}{\Phi}$,
where $g\in\operatorname{GU}\nolimits_{n}(q)$, $e$ is a diagonal element of $\operatorname{GU}\nolimits_{n}(q)$ of order $p$, and $d\in\operatorname{SU}\nolimits_{n}(q)$. The second and third cases occur only in Theorem 4.7, and we have employed that $g\in\langle x,Y,d\rangle$ with $x,\Phi]=[Y,\Phi]=1$ by Lemma 4.6(e). Moreover, we clearly have
(20) $\Phi$ and $^{{[d]}\!}{\Phi}$ belong to $\operatorname{PSU}\nolimits_{n}(q)\langle\Phi\rangle$.
We show next that, if $n_{p}<(q+1)_{p}$, then we also have
(21) $^{{[g]}\!}{[}e]\in\operatorname{PSU}\nolimits_{n}(q),$
where $^{{[g]}\!}{[}e]$ is the first generator listed in Equation (19). In fact, as $[g]$ normalizes $\operatorname{PSU}\nolimits_{n}(q)$ by Equation (2), it is enough to show that $e]\in\operatorname{PSU}\nolimits_{n}(q)$. This is a well known fact, see for instance the proof of Lemma [4, Lemma 4.7(a)], and we provide details: As $n_{p}<(q+1)_{p}$, there exists an element $\omega$ in $\mathbb{F}^{*}_{q^{2}}$ of order $\operatorname{ord}(\omega)=pn_{p}$. In particular, $w^{q+1}=1$ and we may assume that $e=\operatorname{diag}(\mu^{i_{1}},\ldots,\mu^{i_{n}})$ with $\mu=\omega^{n_{p}}$. Then, if we write $n=n_{p}n^{\prime}$ with $(n^{\prime},p)=1$, we have, for any $m\in\mathbb{Z}$, that
$e]=[e\operatorname{diag}(\omega^{m},\ldots,\omega^{m})]\text{ and }\det(e\operatorname{diag}(\omega^{m},\ldots,\omega^{m}))=\omega^{n_{p}(i+n^{\prime}m)},$
where $i=i_{1}+\ldots+i_{n}$. Thus, to obtain that $e]\in\operatorname{PSU}\nolimits_{n}(q)$, it is enough to find $m$ with $i+n^{\prime}m\equiv 0\pmod{p}$, and this is possible because $n^{\prime}$ is coprime to $p$. We conclude from Equations (19), (20), and (21), that, if $n_{p}<(q+1)_{p}$, then
(22) $\widetilde{H}_{n-2}(|\mathcal{A}_{p}(\operatorname{PSU}\nolimits_{n}(q))|;\mathbb{Z})\neq 0\text{ and }\widetilde{H}_{n-1}(|\mathcal{A}_{p}(\operatorname{PSU}\nolimits_{n}(q)\langle\Phi\rangle|;\mathbb{Z})\neq 0,$
under the assumptions of Theorems 4.3 or 4.7, respectively, and where we have employed that $\operatorname{PGU}\nolimits_{n}(q)$ and $\operatorname{PSU}\nolimits_{n}(q)$, as well as its extensions by $\langle\Phi\rangle$, have the same $p$-rank by if $n_{p}<(q+1)_{p}$ by Proposition 2.2.
We now come back to the statement of Theorem B and treat first the low dimensional cases $n=2,3$: For the case $n=2$, we have $\operatorname{PSU}\nolimits_{2}(q)\cong\operatorname{PSL}\nolimits_{2}(q)$, and $p$-extensions of $\operatorname{PSL}\nolimits_{2}(q)$ satisfy $\mathcal{Q}\mathcal{D}_{p}$ by [4, Theorem 3.1], but for the excluded cases $(p,q)=(3,2)$, and $(p,q)=(3,8)$ if $B$ contains field automorphisms. If $n=3$ and $B$ does not contain fields automorphisms, then $\operatorname{PSU}\nolimits_{3}(q)B$ satisfies $\mathcal{Q}\mathcal{D}_{p}$ by [4, Proposition 4.8] but for $3\geq q(q-1)$, i.e., for the excluded case $(p,q)=(3,2)$. If $n=3$ and $B$ contains field automorphisms, then we have $n_{p}\leq 3\leq p$, and $p^{2}\leq(q+1)_{p}$ by [1], so that $n_{p}<(q+1)_{p}$. Thus, in the rest of the proof, we may assume that
(23) $\text{either }n\geq 3\text{ and }n_{p}<(q+1)_{p},\text{ or }n\geq 4\text{ and }n_{p}\geq(q+1)_{p}.$
For conciseness, we explicitly consider the possible generators of $B$: a diagonal automorphism of order $p$, $c_{[g]}$ and a field automorphism of order $p$, $c_{[h]}\Phi^{i}c_{[h^{-1}]}$, where $i$ is co-prime to $p$, and $h,g\in\operatorname{GU}\nolimits_{n}(q)$, see Subsection 2.2. After raising to the appropriate power and conjugating $\operatorname{PSU}\nolimits_{n}(q)$ by $h^{-1}]$, we may assume that,
@@
(i) $B=1$: Then $LB=L$ and, if $n\geq 3$ and $n_{p}<(q+1)_{p}$, we are done by Equation (22), where we exclude the case $(p,q)=(3,2)$. Otherwise, by Equation (23), we have $n\geq 4$ and $n_{p}\geq(q+1)_{p}$, so that $p\mid n$ and, by [4, Lemma 4.7(b)],
@@
where the groups involved have $p$-rank n2n-2. As $n-1\geq 3$ and $(n-1)_{p}=1$, we can apply the case above to the left-hand side.
(ii) $B$ is cyclic generated by a field automorphism of order $p$: If $n\geq 3$ and $n_{p}<(q+1)_{p}$, we are done by Equation (22), where we are excluding the case $(p,q)=(3,8)$. Otherwise, $n\geq 4$ and $n_{p}\geq(q+1)_{p}$, and we use [4, Lemma 4.7(b)] upon the addition of field automorphisms to obtain
@@
where the groups involved have $p$-rank n1n-1. Note that $B$ acts on $\operatorname{PSU}\nolimits_{n-1}(q)$ by Equation (24). As $n-1\geq 3$ and $(n-1)_{p}=1$, we may apply here the previous case $n_{p}<(q+1)_{p}$ of this point (ii) to the left-hand side.
(iii) $B$ is cyclic generated by a diagonal automorphisms of order $p$: Equation (2) gives $\mathcal{A}_{p}(\operatorname{PSU}\nolimits_{n}(q)B)=\mathcal{A}_{p}(\operatorname{PGU}\nolimits_{n}(q))$, and this case follows from Theorem 4.3 and the fact that $\operatorname{m}_{p}(\operatorname{PSU}\nolimits_{n}(q)B)=n-1$ by Proposition 2.2.
(iv) $B$ has $p$-rank 22 and it is generated by a field and a diagonal automorphism of order $p$: By Equation (24), we have $B=\langle c_{[g]},\Phi\rangle$ and, by Equations (19) and (20), to show that the cycle constructed in Theorem 4.7 belongs to $|\mathcal{A}_{p}(\operatorname{PSU}\nolimits_{n}(q)B)|$, it is enough to show that the generators of the form $^{{[g^{\prime}]}\!}{[e]}$ belong to $\operatorname{PSU}\nolimits_{n}(q)\langle c_{[g]}\rangle$, and this holds by Equation (2) again. As $\operatorname{m}_{p}(\operatorname{PSU}\nolimits_{n}(q)B)=n$ by Proposition 2.2, we conclude that $\operatorname{PSU}\nolimits_{n}(q)B$ has $\mathcal{Q}\mathcal{D}_{p}$. ∎